OpenStudy (shayaan_mustafa):

Hi everyone. Has anyone idea how to show polynomial long division using latex?

OpenStudy (zarkon):

they need to install this package http://www.sixthform.info/maths/?p=84

OpenStudy (shayaan_mustafa):

So tell to the admin of Openstudy to install this package.

OpenStudy (unklerhaukus):

\[3V^2-2V-1=(V-1) \overline {\bigg)3V^2-2V-1}\] \[\qquad\qquad\quad3V+1\]\[=(V-1) \overline {\bigg)3V^2-2V-1}\]\[\qquad\quad\qquad3V^2-3V\]\[\qquad\qquad\qquad\qquad V-1\]\[\qquad\qquad\qquad\qquad \underline{V-1}\]\[\qquad\qquad\qquad\qquad 0\]